OpenStudy (anonymous):

theres 15 jelly beans 5 red 5 blue and 5 green what woud the probabiliti that you pick a red one

OpenStudy (anonymous):

none

OpenStudy (anonymous):

1 ou of 15

OpenStudy (mathstudent55):

Since tehre are 15 jelly beans you can pick, but only 5 are red, \( P(red) = \dfrac{number~of~red~jelly ~beans}{total~number~of~jelly~beans} = \dfrac{5}{15} = \dfrac{1}{3} \)